OpenStudy (anonymous):

how do you do 8 * (-6) showing work

OpenStudy (anonymous):

I know it is -48, but is it just 6 X 48 - -48

OpenStudy (anonymous):

I'm not sure what other work you could show besides the associative property. This just means that we can put the parentheses wherever we want and it won't make a difference to the operator. 8 * (-6) = 8 * -6 = -48

OpenStudy (anonymous):

Ok, I have other questions if you are able to help me. I have to show all work for the practice test I am working on.

OpenStudy (anonymous):

Solve 20 -15 /(-5) I have the answer of 23 and I am not sure how it comes out?

OpenStudy (anonymous):

Is it \(\large 20 - \frac{15}{(-5)}\) OR \[\large \frac{ 20-15 }{ (-5) }\]